# Surrogate Models for Vibrational Entropy Based on a Spatial Decomposition

### 21 Feb 2024 v2

@@ 2. Background and Preliminaries
We consider the site potential to be a collection of mappings $V_{\ell}:(\mathbb{R}^{m})^{\mathcal{R}_{\ell}}\rightarrow\mathbb{R}$, which represent the energy distributed to each atomic site. We make the following assumption on regularity and symmetry: $V_{\ell}\in C^{K}\big{(}(\mathbb{R}^{m})^{\mathcal{R}_{\ell}}\big{)}$ for some $K$ and $V_{\ell}$ is homogeneous outside the defect region, namely, $V_{\ell}=V$ and $\mathcal{R}_{\ell}=\mathcal{R}$ for $\ell\in\Lambda\setminus B_{R^{\rm def}}$. Furthermore, $V$ and $\mathcal{R}$ have the following point symmetry: $\mathcal{R}=-\mathcal{R}$, it spans the lattice $\textrm{span}_{\mathbb{Z}}\mathcal{R}=\Lambda$, and $V\big{(}\{-A_{-\rho}\}_{\rho\in\mathcal{R}}\big{)}=V(A)$. We refer to [10, Section 2.3] for a detailed discussion of those assumptions.
@@ 2.3. Site entropy
To establish the aforementioned spatial decomposition, we employ a self-adjoint operator $\textbf{F}_{N}:\mathcal{W}^{\text{per}}_{N}\rightarrow\mathcal{W}^{\text{per}}_{N}$ which acts as $(H^{\text{hom}}_{N})^{-1/2}$. The existence and properties of $\textbf{F}_{N}$ has been previously established in [6, Lemma 2.5].
@@ 2.4. Thermodynamic limit
Theorem 2.1 (Thermodynamic Limit). [8, Theorem 2.1] Let $\bar{u}\in\dot{\mathcal{W}}^{1,2}$ be a stable solution to (2.16). For $N$ sufficiently large, there exists a locally unique solution to (2.1) such that the following estimates hold:
@@
Furthermore, the thermodynamic limit of $\mathcal{S}_{N}$ as $N\to\infty$, denoted by $\mathcal{S}$, exists. The error in approximating $\mathcal{S}(\bar{u})$ with $\mathcal{S}_{N}(\bar{u}_{N})$ satisfies [6, Theorem 2.3]
@@
Additionally, in [6, Lemma 3.2], it is shown that there exist $\underline{\sigma},\bar{\sigma},\varepsilon>0$ and a contour $\mathcal{C}$ encircling $[\underline{\sigma},\bar{\sigma}]$ but not the origin, such that, if some $u_{\infty}\in\dot{\mathcal{W}}^{1,2}$ satisfies $\sigma\big{(}{\rm\textbf{F}}^{*}H(u_{\infty})\textbf{F}\big{)}\cap(-\underline{\sigma},\infty)\subset[2\underline{\sigma},\overline{\sigma}/2]$ for some $0<\underline{\sigma}<\overline{\sigma}$, then for all $u\in B_{\varepsilon}(u_{\infty})$, the operator $\boldsymbol{\rm F}^{*}H(u)\boldsymbol{\rm F}$ remains uniformly bounded above and below and
@@
Following the arguments in the proof of [6, Proposition 5.8], the existence of a similar limit quantity for $\mathcal{S}_{N,\ell}(\bar{u}_{N})$, namely $\mathcal{S}_{\ell}(\bar{u})$ can be deduced. Thus, instead of analyzing the supercell approximation of site entropy $\mathcal{S}_{N,\ell}(u)$, we will give a locality estimate of $\mathcal{S}_{\ell}(u)$ (cf. Theorem 3.1) in the subsequent section.
@@ 3.2. Numerical validation
where $i$ indexes the lattice points, $\mathcal{N}_{i}$ denotes the set of nearest neighbors of the $i$-th site, $u_{i}$ represents the displacement at the $i$-th lattice site, $C_{ij}$ are the coefficients characterizing the interactions between points $i$ and $j$, and $\delta$ is a constant which controls the contribution of the strength of the cubic term. (Qualitatively, this is not a severe restriction of generality, see e.g. [6, Section 6.2]).
@@
It should be noted that the evaluation of site entropy derivatives is computationally demanding. To address this, we adopted an approach to identify the optimal contour and used the sparse nature of the Hessian Jacobians for derivative calculation. The details of this methodology are elaborated in Appendix C. Additionally, all source codes for the numerical tests can be found in our repository  [^1] https://github.com/tinatorabi/ACEntropy.
@@ 4. Surrogate Models
The computation of vibrational entropy, as outlined in (2.7) within the harmonic approximation, involves an initial step of determining the Hessian matrix, demanding an $O(N^{2})$ calculation. Subsequently, diagonalizing this matrix incurs a scaling of $O(N^{3})$ for systems comprising $N$ atoms. To address this computational challenge, Lapointe et al. [31] proposed a surrogate model for assessing harmonic vibrational entropy using a linear-in-descriptor machine learning (LDML) approach. While their derivation of entropy closely aligns with the methodology presented in Appendix B, their work lacks a demonstration of the crucial property of locality in site entropy. This property is pivotal as it enables the fitting of entropy using descriptors of atomic environments.
In this section we connect our locality results to learning a surrogate for the site entropy functional. Specifically, we employ the Atomic Cluster Expansion (ACE) framework, and present a series of numerical results pertaining to the fitting of entropy and its applications in materials science.
@@ 4.1. Atomic Cluster Expansion (ACE)
Accurate energy and force calculations are best achieved through electronic structure techniques like DFT. However, DFT has limitations for larger systems and longer timescales. To overcome these limitations, machine-learned interatomic potentials (MLIPs), trained on DFT data, have become essential in computational materials science due to their high fidelity [1, 2, 4, 9, 38, 39]. This methodology can be directly transferred to learn surrogate models for entropy.
We utilize the linear atomic cluster expansion (ACE) parameterisation [5]. All in-use MLIPs for materials typically express the total energy, $\mathcal{E}_{\text{ML}}$, as an sum over individual site energies. Leveraging the locality of site entropy, as established in Theorem 3.1, the same approach employed for energy modeling can be seamlessly applied.
To that end, we consider $N$ atoms described by their position vectors $y_{j}$. A set $\mathbf{R}:=\{y_{1},\ldots,y_{N}\}\in\mathbb{R}^{3N}$ of $N$ particle positions is called an atomic configuration. Let $y_{ij}=y_{j}-y_{i}$ be the distances between atom $j$ and a reference atom $i$, and let $\mathbf{y}_{i}=\{y_{ij}\}_{j\neq i}$ represent the atomic neighbourhood around atom $i$. The total entropy of a structure of this kind is broken down into site entropies in the ACE model,
@@ 4.2.1. Models setup
Additionally, we present a comparison between prediciting total entropy by the means of a model fitted only to total and another one only to site entropies. Our training set in both cases included 50 rattled configurations of Silicon (rattling parameter $=0.1\,\text{\AA}\times U(0,1)$) including a vacancy with a $2a_{0}\times 2a_{0}\times a_{0}$ supercell. Our test set included 50 rattled configurations of bulk Silicon (rattling parameter $=0.08\,\text{\AA}\times U(0,1)$). During the training for total entropy, we constructed a model with a body order of 3, a maximum polynomial degree of 16, and a cutoff radius of $5.0$Å. For the site entropy, we assembeled a model with a body order of 4, a maximum polynomial degree of 16, and a cutoff radius of $5.0$Å. The fitting results are showcased in Figure 7. Consistent with expectations, the task of fitting site entropies presents greater complexity relative to the fitting of total entropies. For the specific objective of predicting total entropies, it is empirically more precise to utilize a model that has been exclusively trained on total entropy data.
@@ 4.3. Learning curves
Figure 8 and Figure 9 display the log-log plots corresponding to the total and site entropies, respectively. It is noteworthy that upon comparison, we observe that the convergence happens relatively fast for both cases, which is not unexpected due to the simplicity of our training sets. It is also evident that the total entropy values reach convergence earlier when compared to the site entropies. This observation also aligns with expectations, attributed to the inherently lower complexity associated with fitting total entropy. Figure 10 and Figure 11 show the convergence of our fitting with increasing basis size. As the basis size increases, we observe a convergence in both the site and total entropies, along with their respective derivatives, aligning closely with our expectations.
@@
(a) Total entropy
(b) Total entropy derivatives
Figure 10. RMSE Convergence test for total entropy on Si for increasing basis size, fixing the cutoff at 5.5 Å.
(a) Site entropy
(b) Site entropy derivatives
Figure 11. RMSE Convergence test for site entropy on Si for increasing basis size, fixing the cutoff at 5.5 Å.
@@ 5. Conclusion
In this work, we have demonstrated that the total energy can be decomposed into atomic site contributions and have rigorously estimated the locality of site entropy. Our analysis suggests that vibrational entropy can be accurately predicted using a surrogate model for site entropy, which we have developed using machine learning techniques, specifically employing the Atomic Cluster Expansion (ACE) model. Our numerical experiments primarily focus on point defects such as vacancies and interstitials. We have showcased the robustness of our approach in predicting vibrational entropy and the attempt frequency for transition rates governing point defect migration, which is a critical aspect of transition state theory rate approximations.
While the current study emphasizes the effectiveness of our approach in a relatively straightforward context, it also enables a detailed and rigorous examination, albeit within certain limitations. There are numerous potential generalizations that can be explored, including extending the methodology to more intricate material and defect geometries. Additionally, our results pave the way for predicting various entropy-dependent physical quantities, including but not limited to dynamical properties such as diffusion coefficients.
@@ Appendix A Proofs
In this section, we begin by introducing the necessary concepts and models within the primary context of this work. We accomplish this by conducting a comprehensive review of the framework proposed in [10, 13], while also adapting their approaches to align with the specific objectives of our current work.
@@ A.1. Preliminaries
First of all, we introduce the semi-discrete Fourier transform (SDFT)
@@
with =($\mathcal{R}^{\prime}=(\mathcal{R}\cup\{0\})+(\mathcal{R}\cup\{0\})$, see [13, Section 6.2] for more details. Moreover, the stability implies that $c_{0}|k|^{2}\cdot\mathsf{I}\leq\hat{h}(k)\leq c_{1}|k|^{2}\cdot\mathsf{I}$ with the identity matrix $\mathsf{I}$ for all $k\in\mathcal{B}$ [28]. We observe that $|\hat{h}(k)^{-1/2}|\lesssim|k|^{-1}$ as $|k|\rightarrow 0$, hence we can define
@@
Lemma A.1. [6, Lemma 3.1] Let $F:\Lambda\to\mathbb{R}^{m\times m}$ and F be defined by (A.3) and (A.4), respectively. Then, we have: (i) for any $\bm{\rho}\in\mathcal{R}^{j}$, $j\geq 0$, there exists a constant $C$ such that
@@
Lemma A.2. [6, Lemma 3.3 (i)] Let $X$ be a Hilbert space. Let $A\in\mathcal{L}(X,X)$ be a bounded linear operator with range of finite dimension at most $r\in\mathbb{N}$ and $I+A$ is invertible, then there exist $c_{1}(A),\dots,c_{r+1}(A)\in\mathbb{R}$ such that
@@
Since the resolvent $\mathscr{R}_{z}(u)$ exists for all $z\in\mathbb{C},u\in\mathcal{U}$, the inverse on the right hand side exists as well. Additionally, the term $B^{M}:=\mathscr{R}_{z}^{M}\boldsymbol{\rm F}^{*}\big{(}H^{M}(u)-H(u)$ has finite-dimensional range since this is clearly the case for $H^{M}(u)-H(u)$. Thus, according to Lemma A.5 it follows that
@@
for $11\leq j\leq r+$ 1. Using the techniques shown in [6, Eq. (7.13)], one can obtain
@@
Hence, applying [6, Eq. (3.19) and Eq. (3.26)], we have
@@
$\lesssim\sum_{n\in\Lambda}\big{|}DF(n-\alpha)\big{|}\cdot\big{|}DF(n-\beta)\big{|}$
$\lesssim\sum_{n\in\Lambda}|n-\alpha|^{-d}|n-\beta|^{-d},$
where the last inequality follows from Lemma A.1. Applying Lemma  A.1, we have
@@
$\left|\frac{\partial\mathcal{S}_{\ell}(u)}{\partial u_{n}}\right|$ $\leq\frac{1}{2}\frac{1}{2\pi i}\mathrm{Trace}\left|\oint_{\mathcal{C}}\log^{+}(z)\cdot\left[\mathscr{R}_{z}\mathbf{F}^{*}\frac{\partial H(u)}{\partial u_{n}}\mathbf{F}\mathscr{R}_{z}\ \right]_{\ell\ell}{\rm d}z\right|$ (A.16)
$\lesssim\frac{1}{4\pi i}\oint_{\mathcal{C}}\big{|}\log^{+}(z)\big{|}\cdot|n-\ell|^{-2d}\,{\rm d}z$

### 18 Mar 2024 v3

@@ 2. Background and Preliminaries
We consider the site potential to be a collection of mappings $V_{\ell}:(\mathbb{R}^{m})^{\mathcal{R}_{\ell}}\rightarrow\mathbb{R}$, which represent the energy distributed to each atomic site. We make the following assumption on regularity and symmetry: $V_{\ell}\in C^{K}\big{(}(\mathbb{R}^{m})^{\mathcal{R}_{\ell}}\big{)}$ for some $K$ and $V_{\ell}$ is homogeneous outside the defect region, namely, $V_{\ell}=V$ and $\mathcal{R}_{\ell}=\mathcal{R}$ for $\ell\in\Lambda\setminus B_{R^{\rm def}}$. Furthermore, $V$ and $\mathcal{R}$ have the following point symmetry: $\mathcal{R}=-\mathcal{R}$, it spans the lattice $\textrm{span}_{\mathbb{Z}}\mathcal{R}=\Lambda$, and $V\big{(}\{-A_{-\rho}\}_{\rho\in\mathcal{R}}\big{)}=V(A)$. We refer to [13, Section 2.3] for a detailed discussion of those assumptions.
@@ 2.3. Site entropy
To establish the aforementioned spatial decomposition, we employ a self-adjoint operator $\textbf{F}_{N}:\mathcal{W}^{\text{per}}_{N}\rightarrow\mathcal{W}^{\text{per}}_{N}$ which acts as $(H^{\text{hom}}_{N})^{-1/2}$. The existence and properties of $\textbf{F}_{N}$ has been previously established in [7, Lemma 2.5].
@@ 2.4. Thermodynamic limit
Theorem 2.1 (Thermodynamic Limit). [9, Theorem 2.1] Let $\bar{u}\in\dot{\mathcal{W}}^{1,2}$ be a stable solution to (2.16). For $N$ sufficiently large, there exists a locally unique solution to (2.1) such that the following estimates hold:
@@
Furthermore, the thermodynamic limit of $\mathcal{S}_{N}$ as $N\to\infty$, denoted by $\mathcal{S}$, exists. The error in approximating $\mathcal{S}(\bar{u})$ with $\mathcal{S}_{N}(\bar{u}_{N})$ satisfies [7, Theorem 2.3]
@@
Additionally, in [7, Lemma 3.2], it is shown that there exist $\underline{\sigma},\bar{\sigma},\varepsilon>0$ and a contour $\mathcal{C}$ encircling $[\underline{\sigma},\bar{\sigma}]$ but not the origin, such that, if some $u_{\infty}\in\dot{\mathcal{W}}^{1,2}$ satisfies $\sigma\big{(}{\rm\textbf{F}}^{*}H(u_{\infty})\textbf{F}\big{)}\cap(-\underline{\sigma},\infty)\subset[2\underline{\sigma},\overline{\sigma}/2]$ for some $0<\underline{\sigma}<\overline{\sigma}$, then for all $u\in B_{\varepsilon}(u_{\infty})$, the operator $\boldsymbol{\rm F}^{*}H(u)\boldsymbol{\rm F}$ remains uniformly bounded above and below and
@@
Following the arguments in the proof of [7, Proposition 5.8], the existence of a similar limit quantity for $\mathcal{S}_{N,\ell}(\bar{u}_{N})$, namely $\mathcal{S}_{\ell}(\bar{u})$ can be deduced. Thus, instead of analyzing the supercell approximation of site entropy $\mathcal{S}_{N,\ell}(u)$, we will give a locality estimate of $\mathcal{S}_{\ell}(u)$ (cf. Theorem 3.1) in the subsequent section.
@@ 3.2. Numerical validation
where $i$ indexes the lattice points, $\mathcal{N}_{i}$ denotes the set of nearest neighbors of the $i$-th site, $u_{i}$ represents the displacement at the $i$-th lattice site, $C_{ij}$ are the coefficients characterizing the interactions between points $i$ and $j$, and $\delta$ is a constant which controls the contribution of the strength of the cubic term. (Qualitatively, this is not a severe restriction of generality, see e.g. [7, Section 6.2]).
@@
It should be noted that the evaluation of site entropy derivatives is computationally demanding. To address this, we adopted an approach to identify the optimal contour and used the sparse nature of the Hessian Jacobians for derivative calculation. The details of this methodology are elaborated in Appendix C. Additionally, all source codes for the numerical tests can be found in our repository [43].
@@ 4. Surrogate Models
The computation of vibrational entropy, as outlined in (2.7) within the harmonic approximation, involves an initial step of determining the Hessian matrix, a difficult computation, to perform that typically scales as $O(N^{2})$ if performed with finite differences or automatic differentiation, and challenging to implement “by hand”. Subsequently, diagonalizing this matrix incurs a scaling of $O(N^{3})$ for systems comprising $N$ atoms. Exploiting sparsity can only partially reduce this cost. To address this computational challenge, Lapointe et al. [35] proposed a surrogate model for assessing harmonic vibrational entropy using a linear-in-descriptor machine learning (LDML) approach. Their derivation of entropy closely aligns with the methodology presented in Appendix B, and is formally equivalent to our own derivation. The key justification for either approach is our proof of locality of site entropy.
In this section we connect our locality results to learning a surrogate for the site entropy functional. Specifically, we employ the Atomic Cluster Expansion (ACE) framework, and present a series of numerical results pertaining to the fitting of entropy and its applications in materials science, but with a focus on demonstrating the validity of our analysis in the previous sections.
@@ 4.1. Atomic Cluster Expansion (ACE)
Accurate energy and force calculations are best achieved through electronic structure techniques like DFT. However, DFT has limitations for larger systems and longer timescales. To overcome these limitations, machine-learned interatomic potentials (MLIPs), trained on DFT data, have become essential in computational materials science due to their high accuracy and transferrability [1, 3, 5, 11, 49, 41]. This methodology can be directly transferred to learn surrogate models for entropy.
We utilize the linear atomic cluster expansion (ACE) parameterisation [6, 49]. All in-use MLIPs for materials typically express the total energy, $\mathcal{E}_{\text{ML}}$, as an sum over individual site energies. Leveraging the locality of site entropy, as established in Theorem 3.1, the same approach employed for energy modeling can be seamlessly applied.
To that end, we consider $N$ atoms described by their position vectors $y_{j}$. A set $\mathbf{R}:=\{y_{1},\ldots,y_{N}\}\in\mathbb{R}^{3N}$ of $N$ particle positions is called an atomic configuration. Let $y_{ij}=y_{j}-y_{i}$ be the distances between atom $j$ and a reference atom $i$, and let $\mathbf{y}_{i}=\{y_{ij}\}_{j\neq i}$ represent the atomic environment around atom $i$. The total entropy of a structure of this kind is broken down into site entropies in the ACE model,
@@
We will explore the effects of estimating the site entropy parameterization $S^{*}$ from either total entropy, or from site entropies. The latter has the advantage that it increases the amount of available data, but the potential disadvantage that it employs an ad hoc spatial decomposition. (The decomposition used in out proofs may appear natural but in fact there are infinitely many possible alternative decompositions.)
@@ 4.2.1. Models setup
Additionally, we present a comparison between prediciting total entropy by the means of a model fitted only to total and another one only to site entropies. Our training set in both cases included 50 rattled configurations of Silicon (rattling parameter $=0.1\,\text{\AA}\times U(0,1)$) including a vacancy within a $2a_{0}\times 2a_{0}\times a_{0}$ supercell. Our test set included 50 rattled configurations of bulk Silicon (rattling parameter $=0.08\,\text{\AA}\times U(0,1)$). During the training for total entropy, we constructed a model with a body order of 3, a maximum polynomial degree of 14, and a cutoff radius of $5.0$Å. For the site entropy, we assembeled a model with a body order of 3, a maximum polynomial degree of 14, and a cutoff radius of $5.0$Å. The fitting results are showcased in Figure 7. Consistent with expectations, the task of fitting site entropies presents greater complexity relative to the fitting of total entropies. For the specific objective of predicting total entropies, it is empirically more precise to utilize a model that has been exclusively trained on total entropy data. This empirical observation strongly suggests that there exists an alternative spatial decomposition of total entropy that is “easier to fit” (e.g. more local, lower body-order, smoother) than the explicit construction we employed in our analysis.
@@ 4.3. Learning curves
Figure 8 and Figure 9 display the log-log plots corresponding to the total and site entropies, respectively. It is noteworthy that upon comparison, we observe that the convergence happens relatively fast for both cases, which is not unexpected due to the simplicity of our training sets. It is also evident that the total entropy values reach convergence earlier when compared to the site entropies. This observation also aligns with expectations, attributed to the inherently lower complexity associated with fitting total entropy.
To conduct a quantitative analysis on model performance regarding the prediction of total entropies, with fitting results shown in Figure 7, we trained two distinct models: one fitted exclusively against total entropies, and the other against site entropies. Our first objective was to investigate how variations in the number of training configurations influenced the RMSE observed on the test set. We used training sets of varying sizes, containing rattled configurations of Silicon with a vacancy (with a rattling parameter of $0.1\text{\AA}\times U(0,1)$) within a $2a_{0}\times 2a_{0}\times a_{0}$ supercell. We used a test set consisting of 50 rattled bulk Silicon configurations (rattling parameter being $0.08\text{\AA}\times U(0,1)$). The training was done using a body order of 3 and a polynomial degree of 14. The fitting results, corresponding to two differing cut-off radii, $5.0\text{\AA}$ and $7.0\text{\AA}$, are depicted in Figure 9(a). Furthermore, Figure 9(b) showcases the convergence behavior of these models as we increase the basis size, while maintaining a constant cut-off radius of $5.0\text{\AA}$ and using 50 configurations for training. A convergence trend is observed in both models with increasing basis size, in line with our expectations.
@@
(a) Effect of number of configurations
(b) Effect of basis size
Figure 10. RMSE Convergence for total entropy for Si. Figure (A) shows the RMSE error of total entropy evaluated on the test set using two different models. One is solely trained against site entropies (blue) and once is solely trained against total entropies (green) for two different values of cut-off radius, $r_{c}=5.0$  (solid line) and $r_{c}=7.0$Å  (dashed line). Figure (B) showcases the RMSE error of total entropy evaluated on the test set for these two models with a fixed cut off radius of $5.0$  on a dataset containing 50 training configurations for varying basis size.
@@ 5. Conclusion
In this work, we have demonstrated that the total entropy can be decomposed into atomic site contributions and have rigorously estimated the locality of site entropy. Our analysis suggests that vibrational entropy can be accurately predicted using a surrogate model for site entropy, which we have developed using machine learning techniques, specifically employing the Atomic Cluster Expansion (ACE) model. Our numerical experiments primarily focus on point defects such as vacancies and interstitials. We have showcased the robustness of our approach in predicting vibrational entropy and the attempt frequency for transition rates governing point defect migration, which is a critical aspect of transition state theory rate approximations.
While the current study emphasizes the effectiveness of our approach in a relatively straightforward context, it also enables a detailed and rigorous examination, albeit within certain limitations. There are numerous potential generalizations that can be explored, including extending the methodology to more intricate material and defect geometries.
@@ Appendix A Proofs
In this section, we begin by introducing the necessary concepts and models within the primary context of this work. We accomplish this by conducting a comprehensive review of the framework proposed in [13, 18, 23], while also adapting their approaches to align with the specific objectives of our current work.
@@ A.1. Preliminaries
We introduce the semi-discrete Fourier transform (SDFT)
@@
with =($\mathcal{R}^{\prime}=(\mathcal{R}\cup\{0\})+(\mathcal{R}\cup\{0\})$, see [18, Section 6.2] for more details. Moreover, the stability implies that $c_{0}|k|^{2}\cdot\mathsf{I}\leq\hat{h}(k)\leq c_{1}|k|^{2}\cdot\mathsf{I}$ with the identity matrix $\mathsf{I}$ for all $k\in\mathcal{B}$ [32]. We observe that $|\hat{h}(k)^{-1/2}|\lesssim|k|^{-1}$ as $|k|\rightarrow 0$, hence we can define
@@
Lemma A.1. [7, Lemma 3.1] Let $F:\Lambda\to\mathbb{R}^{m\times m}$ and F be defined by (A.3) and (A.4), respectively. Then, we have: (i) for any $\bm{\rho}\in\mathcal{R}^{j}$, $j\geq 0$, there exists a constant $C$ such that
@@
Lemma A.2. [7, Lemma 3.3 (i)] Let $X$ be a Hilbert space. Let $A\in\mathcal{L}(X,X)$ be a bounded linear operator with range of finite dimension at most $r\in\mathbb{N}$ and $I+A$ is invertible, then there exist $c_{1}(A),\dots,c_{r+1}(A)\in\mathbb{R}$ such that
@@
Since the resolvent $\mathscr{R}_{z}(u)$ exists for all $z\in\mathbb{C},u\in\mathcal{U}$, the inverse on the right hand side exists as well. Additionally, the term $B^{M}:=\mathscr{R}_{z}^{M}\boldsymbol{\rm F}^{*}\big{(}H^{M}(u)-H(u)\big{)}{\bf F}$ has finite-dimensional range since this is clearly the case for $H^{M}(u)-H(u)$. Thus, according to Lemma A.5 it follows that
@@
for $11\leq j\leq r+$ 1. Using the techniques shown in [7, Eq. (7.13)], one can obtain
@@
Hence, applying [7, Eq. (3.19) and Eq. (3.26)], we have
@@
$\lesssim\sum_{\ell\in\Lambda}\big{|}DF(\ell-\alpha)\big{|}\cdot\big{|}DF(\ell-\beta)\big{|}$
$\lesssim|n-\alpha|^{-d}|n-\beta|^{-d},$
where we used the fact that $D_{\tau}e_{n}(\ell)$ is non-zero only for $\ell$ in a bounded neighbourhood of $\ell$, and the last inequality follows from Lemma A.1. Applying Lemma  A.1, we have
@@
$\left|\frac{\partial\mathcal{S}_{\ell}(u)}{\partial u_{n}}\right|$ $\leq\frac{1}{2}\frac{1}{2\pi}\mathrm{Trace}\left|\oint_{\mathcal{C}}\log^{+}(z)\cdot\left[\mathscr{R}_{z}\mathbf{F}^{*}\frac{\partial H(u)}{\partial u_{n}}\mathbf{F}\mathscr{R}_{z}\ \right]_{\ell\ell}{\rm d}z\right|$ (A.16)
$\lesssim\frac{1}{4\pi}\oint_{\mathcal{C}}\big{|}\log^{+}(z)\big{|}\cdot|n-\ell|^{-2d}\,{\rm d}z$